OpenStudy (anonymous):

Determine whether the sequence converges or diverges. If it converges, give the limit. 60, -10,5/3 ,5/18 , ...

OpenStudy (cwrw238):

lets see if we have a common ratio here: -10/60 = -1/6, 5/3 / -10 = -5/30 = -1/6 5/18 / 5/3 = -6 ok - are you sure you have the question right?

OpenStudy (cwrw238):

* 1/6

OpenStudy (cwrw238):

is the second term 10 , not -10?

OpenStudy (anonymous):

oh my bad!! it's -10 and -5/18

OpenStudy (cwrw238):

oh ok so r = -1/6 so we have a geometric sequence which oscillates

OpenStudy (anonymous):

so is it converging with a limit of 0?

OpenStudy (cwrw238):

looks like but i'd have to check..

OpenStudy (anonymous):

okay let me know, please

OpenStudy (anonymous):

is that correct?

OpenStudy (cwrw238):

the sum to infinity is a / 1 - r = 60 /( 1 - (-1/6) = 60 / 7/6 = 360/7 = 51.43

OpenStudy (cwrw238):

a GP with -1 < r < 1 converges to a limit

OpenStudy (anonymous):

these are the answer choices: Diverges Converges; 11100 Converges; 72 Converges; 0

OpenStudy (cwrw238):

converges 0 is correct the terms come very close to zero at infinity